# AI-Empowered Integrated Sensing and Communications

### v1

@@ VI-B2 Hybrid Beamforming Design
The vectorization operator is performed on the matrices $\mathbf{F}_{\mathrm{opt}}$, $\mathbf{F}_{\mathrm{RF}}$, and $\mathbf{F}_{\mathrm{BB}}$, resulting in $\tilde{\mathbf{z}}=\text{vec}(\mathbf{F}_{\mathrm{opt}})$, $\tilde{\mathbf{x}}=\text{vec}(\mathbf{F}_{\mathrm{RF}})$, and $\tilde{\mathbf{B}}=\mathbf{F}_{\mathrm{BB}}\otimes\mathbf{I}_{N_{t}}$, where $\otimes$ denotes the Kronecker product. Then the resulting complex-valued quantities are transformed into real-valued ones, denoted as
$\mathbf{z}\triangleq\begin{bmatrix}\mathfrak{R}(\tilde{\mathbf{z}})\\ \mathfrak{I}(\tilde{\mathbf{z}})\end{bmatrix},\mathbf{x}\triangleq\begin{bmatrix}\mathfrak{R}(\tilde{\mathbf{x}})\\ \mathfrak{I}(\tilde{\mathbf{x}})\end{bmatrix},\mathbf{B}\triangleq\begin{bmatrix}\mathfrak{R}(\tilde{\mathbf{B}})&-\mathfrak{I}(\tilde{\mathbf{B}})\\ \mathfrak{I}(\tilde{\mathbf{B}})&\mathfrak{R}(\tilde{\mathbf{B}})\end{bmatrix},$ (31)
where $\mathfrak{R}(\cdot)$ and $\mathfrak{I}(\cdot)$ denote the the real and imaginary parts of a complex-valued vector (matrix), respectively.
Thus, problem (30) can be reformulated as a least-squared problem, denoted as
$\mathbf{x}_{\mathrm{ML}}=\underset{{\mathbf{x}\in\mathcal{S}}}{\operatorname*{\operatorname*{\operatorname*{argmin}}}}\left\|\mathbf{z}-\mathbf{B}\mathbf{x}\right\|^{2},$ (32)
where $\mathcal{S}$ is defined as
$\mathcal{S}\triangleq\{\mathbf{x}:x_{i}+jx_{N_{\mathrm{t}}N_{\mathrm{RF}}+i}=\tilde{x}_{i},|\tilde{x}_{i}|=1,i=1,\ldots,N_{\mathrm{t}}N_{\mathrm{RF}}\}.$
This constrained least-squares problem can be solved using DetNet. Specifically, the output of $l$th layer of the neural network ($\mathbf{x}_{l}$) can be expressed as
[…]
where $\delta_{\ell}$ is the stepsize and $\Pi_{\ell}$ nonlinear projection operator, which are learned by proposed DNN architecture (see architecture details in [148]). The employed loss function is
$\mathcal{L}=\sum_{\ell=1}^{L}\log{\ell}\left\|\mathbf{z}-\mathbf{B}\mathbf{x}_{\ell}\right\|^{2},$ (34)
where $L$ is the number of layers of the unrolled network.
Once offline training is completed, the well-trained neural network is used to perform online HBF design. Then, $\mathbf{F}_{{\mathrm{RF}}}$ can be reconstructed from $\mathbf{x}_{L}$. Define $\tilde{\mathbf{H}}=\mathbf{H}\mathbf{F}_{{\mathrm{RF}}}$ and $\mathbf{Q}=\mathbf{F}^{H}_{{\mathrm{RF}}}\mathbf{F}_{{\mathrm{RF}}}$. The digital precoder design problem can be expressed as [148]
[…]
which has the well-known water-filling solution:
$\mathbf{F}_{{\mathrm{BB}}}=\mathbf{Q}^{{-\frac{1}{2}}}\tilde{\mathbf{U}}\tilde{\bm{\Gamma}},$ (36)
where the columns of $\tilde{\mathbf{U}}$ are taken from the right singular vectors corresponding to the $N_{s}$ largest singular values of $\tilde{\mathbf{H}}\mathbf{Q}^{{-\frac{1}{2}}}$, and $\tilde{\bm{\Gamma}}$ is a diagonal matrix whose elements are de- fined by the power allocated to the $N_{s}$ data streams.
VI-B3 Numerical Results
In this section, the performance of the rolled neural network tailored to the HBF design is tested. The system parameters are set as $N_{t}=32$, $N_{s}=N_{\mathrm{RF}}=2$. The channel $\mathbf{H}$ is modeled as Rayleigh fading channel. The sizes of datasets for training and testing are 500 and 100. For training phase, Adam optimizer with a learning rate of 0.01 is used. In Figs. 10 and 10, the HBF design using unrolled neural network is named as ‘unrolled HBF’, and unconstrained fully DBF as ‘optimal DBF’.
The spectral efficiency versus the SNR for $N_{r}=16$ and $N_{r}=32$ are shown in Fig. 10. It is observed that for both cases, the spectral efficiency gaps between optimal DBF and unrolled HBF for $\text{SNR}=-10\,\text{dB}$ is smaller than that for $\text{SNR}=15\,\text{dB}$. In addition, the spectral efficiency gap between optimal DBF and unrolled HBF for $N_{r}=32$ is slightly smaller than that for $N_{r}=16$ across the whole SNR region. This implies that the spectral efficiency performance of unrolled HBF scheme is close to that of optimal BF when the number of receiving antennas is large and the SNR is in low region.
Next, the spectral efficiency performance versus the number of receiving antennas is tested in the setting where the SNR is $10$ dB. As shown in Fig. 10, as the number of receiving antennas increases, the spectral efficiency gap between optimal BF and unrolled HBF gets smaller, which means that the unrolled HBF scheme in our context performs well when the number of receiving antennas is large.
1010-10- 1088-8- 866-6- 644-4- 422-2- 20022224444666688881010101012121212141414141616161655551010101015151515SNR (dB)Spectral Efficiency (bits/s/Hz)Optimal DBF, $N_{r}=32$Unrolled HBF, $N_{r}=32$Optimal DBF, $N_{r}=16$Unrolled HBF, $N_{r}=16$
Figure 9: Spectral efficiency versus SNR for $N_{t}=32$, $N_{s}=N_{\mathrm{RF}}=2$, $N_{r}=16$ and $N_{r}=32$.
[…]
Figure 10: Spectral efficiency versus the number of receiving antennas for $N_{t}=32$, $N_{s}=N_{\mathrm{RF}}=2$, SNR is $10$ dB.
VII Future Research Directions
@@
The range of problems in ISAC, which have been shown promise to benefit from AI is wide. Here, we aim to identify several promising directions for future research in the area of AI-empowered ISAC. This list is not intended to be complete, but it focuses more on ISAC open problems, particularly in waveform and beamforming design.
VII-A End-to-End ISAC
@@
While AI-based constellation design for interference mitigation in communications has been under consideration recently [126, 123, 13], there is currently no work for ISAC. This is due to the unique challenges posed by ISAC, where a unified waveform is used for both communication and sensing tasks, leading to overlapping signal components in temporal, frequency, and spatial domains. Future directions in end-to-end learning-based ISAC can focus on addressing this problem. In addition, addressing practical challenges, such as minimizing symbol error rate in the presences of interference in various forms, including self-interference, inter-cell interference, and inter-user interference. Further, extending these methods to multi-user and multi-target scenarios, while enhancing spectral efficiency and ensuring scalability, will be critical to realizing the full potential of ISAC systems in practical deployments.
@@
While current approaches [130, 131, 132] have demonstrated promising path in jointly optimizing sensing and communication tasks, their performance, particularly in realistic scenarios with interference, remains unknown. These works also ignore interference in its various forms including self-interference [152, 153, 154], inter-cell interference[155], and inter-user interference [156], among others. Future research should explore advanced techniques to integrate interference mitigation into end-to-end learning frameworks for both single-carrier and multi-carrier methods, including OFDM and OTFS. Problem of interference reduction involves using end-to-end learning to develop new, flexible constellations that are specifically designed to be aware of self-interference. To achieve this, the research will leverage autoencoder-based techniques for self-interference cancellation, which is a promising approach since interference can be identified at the transmitter. To address this challenge, this task will utilize unsupervised learning techniques to develop adaptive constellations for ISAC.
VII-B Transformer Neural Networks for ISAC Design
@@
AI techniques can extract 3D features from multimodal inputs such as vision, radar, LiDAR, and position data. Transformer networks [159], the foundation of LLMs like ChatGPT and Gemini, can fuse these features to learn ISAC waveforms and beamformers. Their masked multi-head attention preserves structural dependencies, while customized masking [160] enables embedding of sensing and communication-specific side information. These capabilities make transformers well-suited for addressing complex ISAC optimization problems with improved accuracy and reduced data requirements.
VII-C Digital Twin-aided ISAC System Designs
@@
In the context of ISAC, digital-twin-aided architectures are particularly promising for dynamic and distributed scenarios. By collecting spatial and temporal information from distributed access points and sensors, digital twins can construct 3D models of the surrounding environment, capturing positions, orientations, and shapes of objects such as scatterers, targets, and users. These digital replicas enable simulation of communication and sensing channels, which can be used to design ISAC transmission strategies and train AI models for unified waveform and beamformer design.

### v2

@@ VI-C2 Sensing
in which, $\hat{T}$ represents the DNN estimation, and $T$ denotes the actual presence of the target.
VI-C3 ISAC
In a setting where separate signals are sent for communication and sensing, the above optimizations are straightforward and usually result in QAM-like constellations for communication and constant-modulus modulation, such as PSK, for sensing. However, in an ISAC system where the same signal is used for both sensing and communication—as in our case in Fig. 12—finding an appropriate constellation becomes more challenging, as it must take the competing requirements of both tasks into account.
With the goal of minimizing the BER and maximizing $P_{D}$ using a single signal $x$ in the ISAC system, the overall loss function is formulated as the weighted sum of the communication and radar loss functions, given by
$\mathcal{L}_{\rm ISAC}=\eta\mathcal{L}_{\rm radar}+\bar{\eta}\mathcal{L}_{\rm comm},$ (43)
where $0\leq\eta\leq 1$ is a hyper-parameter that defines the relative importance of communication versus radar. The values of $\eta$ close to 1 correspond to radar-centric scenarios, while values close to 0 correspond to communication-centric scenarios. The loss functions $\mathcal{L}_{\rm comm}$ and $\mathcal{L}_{\rm sens}$ are defined in (41) and (42).
After defining the network structure, we simultaneously learn the neural network models $\Psi_{\rm E}$, $\Psi_{\rm comm}$, and $\Psi_{\rm radar}$ through appropriate training. These correspond to the encoder, communication decoder, and radar detector, respectively. Table VII summarizes the number of neurons in the hidden layers and the activation functions used in each neural network. It can be seen that all three models share the same structure for their hidden layers, but their input and output layers differ. We use the ReLU activation function between hidden layers.
TABLE VII: Neural networks architecture
Model Input Hidden Layers Output
$\Psi_{\rm E}$ $K$ $(16,32,16)$ 22 (linear)
$\Psi_{\rm comm}$ 22 $(16,32,16)$ $K$ (softmax)
$\Psi_{\rm radar}$ 22 $(16,32,16)$ 11 (sigmoid)
Figure 13 shows examples of learning-based constellations. To obtain these, we set the message length to $K=5$, resulting in a constellation size of $M=32$. The model is trained using the Adam optimizer with a learning rate of 0.001 for 500 epochs, with $10^{6}$ samples per epoch and a batch size of $10^{5}$.
As illustrated in Fig. 13(a), the NN-based algorithm generates novel constellations by varying the trade-off parameter $\eta$, which balances sensing and communication performance. Specifically, the figure shows constellations corresponding to $\eta=0.9$ (sensing-centric), $\eta=0.7$ (balanced), and $\eta=0.05$ (communication-centric). It is seen that, PSK-like structures are effective in sensing-centric scenarios, QAM-like structures emerge in communication-centric cases, and multi-ring constellations are optimal when balancing both objectives. In Table VIII, the SER, $P_{D}$, and $P_{\rm FA}$ of the proposed designs are provided. Threshold chosen such that $P_{fa}$ is as small as about $10^{-3}$. For comparison with traditional constellations, we have also included the performance of 32-QAM and 32-PSK in this table.
Figure 13: Comparison of various 32-point constellations designed using neural network for different ISAC scenarios. The value of the trade-off parameter $\eta$ is indicated in each subfigure.
TABLE VIII: Performance metrics for learning-based constellations with different $\eta$ and traditional constellations
Constellation Type SER $P_{d}$ $P_{fa}$
Learning-based $\eta=0.05$ $10^{-1.38}$ 0.157 0.045
$\eta=0.70$ $10^{-1.19}$ 0.570 0.0067
$\eta=0.90$ $10^{-0.46}$ 0.935 0.0088
Traditional 32-QAM $10^{-1.39}$ 0.170 0.0240
32-PSK $10^{-0.49}$ 0.935 0.0085
Other performance metrics may also be used to compare these constellations. In particular, the MMSE of each constellation can be evaluated using [197, Eq. 24]. Furthermore, the CRLB performance of the proposed ML-based constellation represents another important evaluation criterion. The evaluation of the CRLB for phase and frequency estimation in classical modulation schemes has been studied in [198, 199, 200, 201]. Since evaluating the exact CRLB is mathematically involved, the modified CRLB (MCRB) was proposed to provide a more tractable bound [199]. The MCRBs of phase ($\phi$) and frequency ($\nu$) for PSK and QAM modulated signals are given by [199, 201, 200] $\mathrm{MCRB}_{\phi}=\frac{N_{0}}{2E_{s}}\cdot\frac{1}{L}$ and $\mathrm{MCRB}_{\nu}=\frac{N_{0}}{2E_{s}}\cdot\frac{3}{\pi^{2}L(L^{2}-1)}$ where $L$ is the number of samples used for estimation, and $\frac{E_{s}}{N_{0}}$ is the symbol energy-to-noise ratio. The MCRB provides a looser but simpler bound that is accurate at high SNRs and identical for PSK and QAM constellations.
At low SNRs, closed-form CRLBs have been derived only for certain constellations such as BPSK and QPSK [200]. The MCRB becomes less accurate in this regime, and the exact CRLB strongly depends on the modulation type. Since QAM and PSK share similar MCRBs, it is reasonable to expect comparable results for learning-based constellations. However, due to their irregular and asymmetric structures, evaluating low-SNR CRLBs for ML-based constellations remains a challenging, yet interesting open problem.
Finally, range accuracy, can also depend on the constellation, because it is tied to how well the received signal allows estimation of parameters such as delay or phase under noise. This relationship is captured by the CRLB for time-delay (or phase) estimation, given by $\mathrm{CRB}_{\tau}=\frac{1}{8\pi^{2}\beta^{2}\,\mathrm{SNR}}$ [29, Eq. 62] where $\beta$ is the effective root-mean-square bandwidth of the signal. The effective $\beta$ can change slightly depending on the symbol energy distribution, the instantaneous power envelope, and the phase structure of the constellation, and is mostly influenced by pulse shaping.
VII Lessons Learned, Limitations & Future Research Directions
The range of ISAC problems that have shown promise in benefiting from AI is wide. In this section, we first summarize key lessons learned and then discuss directions for future research.
VII-A Lessons Learned
A key takeaway from this tutorial is that ISAC problems are typically trade-off problems—meaning they involve the simultaneous optimization of multiple, often conflicting, objectives. Consequently, regardless of the solution approach—whether traditional multi-objective optimization or emerging learning-based methods—the results often lead to Pareto-optimal solutions. An example of such solutions is one that balances communication throughput and sensing accuracy, where improving one objective typically comes at the cost of the other.
AI-empowered solutions have both advantages and limitations when applied to ISAC systems. We discuss these below, noting that there is currently little guidance on when each approach is most suitable.
VII-A1 When to Use Each AI Model?
In this paper, we have discussed various AI models. Below is a brief summary of when each model is most appropriate for ISAC applications.
  • Single-Task Learning: Designed to optimize performance for one specific task, without leveraging shared features from related tasks. For example, training a model solely for target detection in ISAC.
  • Multi-Task Learning: Suited for two or more related tasks that can share features, e.g., joint channel estimation and symbol detection, or detecting pedestrians and predicting the road’s curvature in a self-driving car model.
  • Multi-Objective Learning: Optimizes multiple, and often competing, objectives in a single task, e.g., a car model trained solely to detect pedestrians (one task) but has two objectives: to maximize accuracy and to minimize inference time.
  • Model-Based Learning: Best for problems with known physical models but practical uncertainties, e.g., hybrid beamforming with impairments or waveform design combining analytical and learned components.
  • Deep Learning: Fits high-dimensional, non-linear problems lacking tractable models, e.g., end-to-end waveform design or complex resource allocation.
  • End-to-End Learning: Trains a model to directly map raw inputs to final outputs, jointly optimizing all intermediate processing stages, e.g., learning ISAC waveforms and detection algorithms in one unified framework without separately optimizing each block.
VII-A2 Benefits of AI-Empowered ISAC Systems
We have discussed several advantages of AI-based ISAC over classical optimization approaches throughout the paper. These advantages can be classified as:
  • Reducing Computational Complexity: Many AI-based ISAC designs belong to this category. For example, as demonstrated in Case Study I, lightweight unsupervised AI models can largely reduce the computational complexity of classical optimization techniques while maintaining a negligible performance or accuracy gap. This gap can be further minimized by increasing dataset size and fine-tuning the model architecture, complexity, and generalization capability through techniques such as regularization and cross-validation.
  • Solving Non-Convex Problems: From Case Study II, we learned that algorithm unfolding is particularly effective for solving non-convex problems in communications and ISAC, as the trained unfolded model can better avoid poor local minima and often achieves improved suboptimal solutions. While the per-layer computations mirror those of the original algorithm, efficiency gains arise from requiring significantly fewer layers than the number of algorithm iterations.
  • Improving Accuracy: From Case Study III, we saw that autoencoder-based constellation design can flexibly balance the sensing–communication trade-off and produce novel, previously unseen solutions that outperform both traditional constellations (e.g., QAM) and probabilistic-shaped constellations in terms of sensing accuracy and communication SER.
  • Simplifying System Design: Some DL models, such as the end-to-end learning, offer a powerful and streamlined alternative to traditional block-by-block system design. By optimizing multiple components, e.g., channel coding, modulation, and beamforming, in a single framework, end-to-end learning reduces system design complexity compared to traditional approaches. This approach could advance the development of AI-native air interfaces, a long-term objective in AI-driven communication systems.
  • Enhancing Robustness: Learning-based approaches can enhance robustness in ISAC systems, particularly when traditional model-based methods struggle with model mismatches or hardware impairments. Such mismatches may arise from imperfect or quantized CSI, rapidly changing channel conditions, hardware limitations, or the inherent complexity of ISAC operations. By leveraging real-time data, AI models can adapt to these uncertainties and design robust waveforms and beamformers compared to convex optimization [19], and demonstrates superior error-rate performance under hardware impairments [20].
VII-A3 Limitations of AI-Based ISAC Systems
Despite its potential benefits, AI-based ISAC system designs have several limitations. These include
  • In supervised ISAC, collecting and labeling large datasets for ISAC systems—especially in dynamic wireless environments—can be challenging and costly. Fortunately, as noted earlier, real-world datasets are becoming increasingly available [38].
  • AI-aided ISAC models trained on specific datasets may not generalize well to new signal characteristics, dynamic propagation conditions, or unseen noise and interference. This limits their robustness in real environments, where signals are constantly affected by multipath fading and co-channel interference.
  • Deep learning models can be hard to interpret due to their complexity and lack of transparency. Such AI models can inherit biases of their training data, leading to unfair or discriminatory outcomes of the end-to-end ISAC systems.
  • Learning-based ISAC waveforms and beamformers may also suffer from generalization limitations. Both sensing and communication performance metrics degrade when the CSI, which is used to train the model, becomes aged/outdated or when the system topology changes. In such cases, the models must be retrained by considering the latest CSI and updating the system topologies.
  • AI-based ISAC may raise privacy and security concerns when processing sensitive data, such as user location, and are vulnerable to attacks such as adversarial examples that can manipulate inputs and disrupt system behavior.
It is worth noting, however, that the last three limitations are currently paid increasingly much attention in several research communities and there are good promisees to overcome them, or develop guarantees on performance, generalizability, and privacy. These are, however, outside of the scope of this tutorial.
VII-B Future Research Directions
@@
Here, we identify several directions for future research in the area of AI-empowered ISAC. This list is not intended to be complete, but it focuses more on ISAC open problems, particularly in waveform and beamforming design.
VII-B1 End-to-End ISAC
@@
While AI-based constellation design for interference mitigation in both communications and radar has been studied recently [164, 161, 32], there is currently no work for ISAC. Future directions in end-to-end learning-based ISAC can focus on addressing this problem, e.g., by extending Case Study III to a multi-cell scenario. In addition, addressing practical challenges, such as minimizing symbol error rate in the presences of interference in various forms, including self-interference, inter-cell interference, and inter-user interference. Further, extending these methods to multi-user and multi-target scenarios, while enhancing spectral efficiency and ensuring scalability, will be critical to realizing the full potential of ISAC systems in practical deployments.
@@
While current approaches [20, 168, 169] have demonstrated promising path in jointly optimizing sensing and communication tasks, their performance, particularly in realistic scenarios with interference, remains unknown. These works also ignore interference in its various forms including self-interference [202, 203], inter-cell interference[204], and inter-user interference [205], among others. Future research should explore advanced techniques to integrate interference mitigation into end-to-end learning frameworks for both single-carrier and multi-carrier methods, including OFDM and OTFS. Problem of interference reduction involves using end-to-end learning to develop new, flexible constellations that are specifically designed to be aware of self-interference. To this end, the research could leverage autoencoder-based techniques for self-interference cancellation, which is a promising since interference can be identified at the transmitter. In addition, a deeper analysis of the low-SNR performance of ML-based constellations, including their CRLB for phase, frequency, and range estimation, is needed to clarify their advantages and limitations. Further, ML-based constellations could potentially be designed to achieve a desired CRLB bound.
VII-B2 Transformer Neural Networks for ISAC
@@
AI techniques can extract 3D features from multimodal inputs (vision, radar, LiDAR, position). Transformer networks [208], the foundation of large language models like ChatGPT and Gemini, can fuse these features to learn ISAC waveforms and beamformers. Unique features of transformers such as long-range dependency modeling, self-attention, parallel processing, and positional encoding can be leveraged for enhancing ISAC designs over the existing deep learning models.
Transformers offer strong potential for ISAC systems by jointly learning sensing and communication tasks through multi-head self-attention. The global self-attention mechanisms offered by generative transformers can be leveraged to dynamically aggregate information across distributed access points (APs), users, and targets without relying on any predefined graphical structures. They enable each network entity (APs, user, and target) to attend to every other entity, thereby capturing full-order dependencies and modeling global cooperation across the distributed ISAC-enabled entities. Their ability to model global dependencies across multimodal data enables efficient unified waveform and beamformer design that can outperform traditional approaches based on convolutional neural networks and graph neural networks (GNN), in dynamic propagation environments envisioned for 6G. Unlike models relying on local or predefined structures, transformers dynamically assign attention weights to all features, allowing task-oriented focus for channel estimation, target tracking, and signal processing. Masked multi-head attention and customized embeddings [209, 210] further integrate sensing and communication side information, while hybrid transformer–GNN models [211, 212] enhance multi-user coordination. Despite challenges such as high computational cost, long-sequence handling, and interpretability, transformers enable faster, parallel optimization of ISAC waveforms and beamforming with superior adaptability in complex environments.
VII-B3 Digital Twin-Aided ISAC Systems
@@
In the context of ISAC, digital-twin-aided architectures are particularly promising for dynamic and distributed scenarios. By collecting spatial and temporal information from distributed APs and sensors, digital twins can construct 3D models of the surrounding environment, capturing positions, orientations, and shapes of objects such as scatterers, targets, and users. These digital replicas enable simulation of communication and sensing channels, which can be used to design ISAC transmission strategies and train AI models for unified waveform and beamformer design.
@@
VII-B4 Learning-Based ISAC in FR3
Most existing ISAC designs focus on sub-6 GHz or mmWave frequencies. However, the upper mid-band spectrum spanning 7–24 GHz, also known as frequency range 3 (FR3) [214, 215], is emerging as a key enabler for 6G ISAC. It offers the potential to bridge the wide-area coverage advantages of sub-6 GHz (FR1) with the high-capacity benefits of mmWave (FR2). However, directly applying learning-based algorithms developed for FR1 or FR2 to FR3 ISAC may not be feasible due to the distinct propagation characteristics, antenna configurations, and temporal dynamics of the FR3 band [215]. FR3 channels exhibit moderate sparsity, higher angular resolution, and mid-range Doppler variations, which require tailored feature representations and model architectures. Moreover, as large-scale arrays in FR3 approach the near-field regime, traditional far-field assumptions—central to many FR1 and FR2 algorithms—break down, necessitating range–angle joint modeling and beam focusing rather than simple steering. To effectively operate in FR3, learning algorithms must integrate frequency-aware normalization, domain adaptation, and physics-informed priors, ensuring that they can leverage FR3’s unique sensing–communication trade-offs while maintaining robustness to moderate blockage, beam misalignment, and frequency-dependent scaling effects.